# A critical assessment of the role of topology on protein thermal stability

### v1

@@ Abstract
For several decades, experimental and computational studies have been used to investigate the potential functional role of knots in protein structures. A property that has attracted considerable attention is thermal stability, i.e., the extent to which a protein retains its native conformation and biological activity at high temperatures, without undergoing denaturation or aggregation. Thermal stability is quantified by the melting temperature , an equilibrium property that corresponds to the peak of heat capacity in differential scanning calorimetry experiments. Since protein energy does not depend on the topological state, equilibrium properties such as should not depend on the topology, provided that the exploration of the equilibrium space is ergodic. However, experimental and computational results reported in the literature provide conflicting views on this problem, with some studies reporting an enhancement of thermal stability for certain knotted proteins. Here, we use extensive Monte Carlo simulations of a simple C-alpha protein model of protein YibK to show in a comprehensive manner that does not depend on the topological state of the protein.
@@ 1 Introduction
A knotted protein is a protein whose polypeptide chain is self-tied, forming a physical (i.e., open) knot in its native structure. Although the first knotted protein was reported in 1977 richardson1977, it was not until 2000 that Taylor’s computational methodology (which enables a systematic determination of the topological state of open polymer chains) taylor2000 brought these complex proteins into the spotlight. Knotted proteins make up 1% of all protein entries (225 K) in the Protein Data Bank (PDB) Sulkowska2018, and are statistically rare compared to compact random loops of the same size Grosberg2006. All three domains of life contain knotted proteins Virnau2006, including Archaea, a group of single-celled organisms believed to have inhabited the planet for around 3.8 billion years. A recent survey that analyzed 700K knotted proteins predicted by the computer artificial intelligence system AlphaFold (AF) AlphaFold, and 42 proteomes (16 Bacteria, 9 Fungi, 4 Plant, and 13 Animal), reported that the percentage of knotted proteins in any given proteome is around 0.4% Sulkowska2024. The minimum number of crossings in all possible planar projections of a given chain is an incomplete topological invariant used to classify both topological knots (which are closed curves in space) and protein knots. The (or trefoil) knots, a class that features three crossings, is by far the most prevalent protein knot, followed by the and knots, with four and five crossings, respectively Sulkowska2018, Sulkowska2024. The most complex protein knot has seven crossings (it is a knot) and was predicted by AF AlphaFold. It is an 89-residue-long bacterial protein (termed Q9PR55) whose atomic structure was recently experimentally resolved, validating AF’s predictions Danny7_1.
@@
In the context of structural biology, the design principle ‘form follows function’ means that the structure of a protein is adapted to successfully perform its specific task. It highlights the fact that protein structure is not arbitrary, being adjusted by the specificities of its role, which ultimately contributes to the organism’s survival. Additionally, the fact that knots in proteins have been conserved throughout evolution Joanna2012 (and, in many cases, have likely evolved from knotted precursor structures in organisms dating back billions of years) suggests that they may confer some advantage to their carriers. Thus, it is not surprising that a large body of theoretical and experimental work has been devoted to determining the potential functional roles of knots in proteins, alongside establishing how they fold into their intricate native structures. While significant advancements have been made regarding folding and knotting mechanisms (reviewed in Tubiana2024), a consensus has not been reached concerning the function of protein knots (reviewed in Jackson2020, Tubiana2024, Danny2023). Knots can potentially influence proteins at both the physical and structural levels, as well as at the level of their biological function. A recent study found that the knot in protein UCHL-1 is critical for the preservation of secondary structure Sara2024. Previous studies have reported that a knotted structure helps shape and form the binding site of enzymes Nureki2002, Nureki2004, Jacobs2002. Additionally, one study has shown that the knotted region is directly involved in enzymatic activity. Indeed, for the protein TrmD (an example of an important class of enzymes, the SPOUT and SAM synthases), it was shown that the knot regulates catalysis, being critical to protein function Ya-Ming2016. Examples of physical properties are thermal stability, which measures a protein’s resistance to melting and is quantified by the melting temperature; kinetic stability, which measures the protein’s resistance to unfolding at a certain temperature and is quantified by the unfolding rate; and mechanical stability, a measure of the protein’s ability to withstand physical forces and deformations without unfolding or losing its native structure. Simulation studies using models with different levels of resolution suggested that knots in proteins enhance kinetic stability Soler2014, Joanna2008, while an enhancement of mechanical stability has been reported for several model systems based on in vitro experiments Danny2020, Danny7_1(reviewed in Danny2023) and computational studies Joanna2008, Xu2018. Some years ago, Monte Carlo simulations of a simple lattice model were used to find that a knotted topology does not affect thermal stability Soler2013. However, this is in contradiction with a recent experimental study that reported a significant enhancement in thermal stability for a protein featuring a deep trefoil knot Danny2019a. The purpose of the present work is to revisit this question. Our results, based on extensive Monte Carlo simulations of a simple off-lattice model, strongly support the view that knots play no role in enhancing a protein’s thermal stability or any other thermodynamic equilibrium property.
@@ 2.1 The Cα Gō model
Proteins are represented by a simple Cα model (Figure 1). Accordingly, residues are reduced to hard spherical beads of uniform size, centered on the Cα atoms. Consecutive Cα atoms are connected by rigid sticks representing pseudobonds on the amide planes. We adopt a radius of 1.7 Å for the beads, which is the van der Waals radius of Cα atoms Tsai1999. For the length of each stick, we adopt the distance between the Cα atoms of the respective bonded residues in the protein’s native conformation, these being approximately 2.9 Å for cis bonds and 3.8 Å for trans bonds. Two non-bonded residues are said to be in contact in the native conformation if the smallest distance between any two heavy atoms, one belonging to each residue, is 4.5 Å, this cut-off being chosen because it is slightly larger than twice the average van der Waals radius of heavy atoms in proteins.
To model protein energetics, we consider the native-centric Gō potential Go. Accordingly, the total energy of a conformation defined by bead coordinates is given by
@@ 2.2 Replica-exchange Monte Carlo simulations
The conformational space of the C-alpha model is explored with Metropolis Monte Carlo (MC) Metropolis(1953) by means of a move set that comprises crankshaft and pivot moves. All simulations start from an unfolded (and unknotted) conformation obtained from a high-temperature simulation, and folding progress is monitored using the fraction of native contacts, . To sample equilibrium distributions, we use Monte Carlo replica-exchange (MC-RE) Sugita(1999).
To explore the conformational space, two sampling schemes can be deployed. A canonical one in which the move set preserves the linear topology of the chain, and another one that breaks the linear topology of the chain by allowing it to cross itself Especial2022, while preserving the sequential bonding of the protein chain and excluded volume among its units. We have previously shown, using simple Cα protein models (of unknotted proteins Fn-III and m, shallowly knotted protein MJ0366, and deeply knotted protein Rds3p), that there is no difference in equilibrium properties calculated with a move set that preserves linear topology and one that does not Especial2022. However, the move set that breaks the linear topology of the chain is extraordinarily advantageous in the case of deeply knotted proteins like the ones studied here, since it provides correct equilibrium results at a much lower computational cost Especial2022. Thus, unless otherwise stated, we will deploy the sampling method that does not preserve the linear topology of the chain.
The weighted histogram analysis method (WHAM) Chodera(2007) is used to analyze data from the MC-RE simulations and produce maximum likelihood estimates of the density of states, from which expected values for thermodynamic properties are calculated as functions of temperature. An example is the heat capacity, , defined in reduced units as . The melting temperature, , is determined as the temperature at which the peaks. We measure the cooperativity degree of the transition by the ratio of the full width at half maximum (FWHM) of the peak to the melting temperature, and the half-width of the potential well, , is adjusted to obtain a simulated FWHM ratio between 4 and 5% Larriva2010, Fernandez2021. This criterion has been successfully used in previous simulations employing a similar potential and sampling faisca2016, faisca2019. WHAM is additionally used to project the density of states along to obtain free energy profiles at a specific temperature.
@@ 2.3 Molecular Dynamics simulations
The MD simulations were performed using GROMACS 2024.3 abraham2015, with the AMBER 14SB force fields maier2015, and the TIP3P water model jorgensen1983, neria1996. The system consisted of a protein, 8808 water molecules, and 4 Cl- ions to equilibrate the overall system charges. We performed a two-step minimization procedure using the steepest descent method and an initialization protocol with two additional steps, in which temperature (NVT) and pressure (NPT) baths are turned on and equilibrated for 100 ps segments, with an integrator step of 1 fs. The nonbonded interactions were computed using a single cutoff of 1.0 nm, with the neighbor list updated every 20 steps. Van der Waals interactions were truncated beyond this cutoff, while long-range electrostatics were calculated using the Particle-Mesh Ewald (PME) method darden1993, with a Verlet cutoff of 1.0 nm and a Fourier grid spacing of 0.125 nm. Bond lengths involving hydrogen atoms in solute and water molecules were constrained using the LINCS hess2008b and SETTLE miyamoto1992 algorithms, respectively. In the NVT ensemble, the v-rescale thermostat bussi2007 was employed to maintain a temperature of 310 K (with a coupling constant of 0.1 ps). In the NPT ensemble, the v-rescale thermostat bussi2007 (coupling constant of 1.0 ps) was used in combination with the Parrinello-Rahman barostat parrinello1981 (coupling constant of 2.0 ps and isothermal compressibility of 4.510-5 bar-1). For the MD relaxation step, we run for 100 ns with the integration step adjusted to 2 fs.
@@ 2.4 Topological state
The topological state (knotted or unknotted) of a sampled conformation is determined using the Koniaris-Muthukumar-Taylor (KMT) algorithm taylor2000.
@@ 3.1 Model systems
Proteins YbeA (from Escherichia coli, PDB ID: 1ns5) and YibK (from Haemophilus influenzae, PDB ID: 1mxi and PDB ID:1j85) are homodimeric enzymes that belong to the \-knot superfamily of MTases. In this work, we consider the monomer identified as chain A in the corresponding PDB files. The native structure of both proteins features a trefoil knot. In a knotted protein, the region of the polypeptide chain encompassing the knot is termed the knotted core. Protein knots are classified as deep or shallow, depending on the number of residues that must be removed from one of the termini to unknot the chain. If the latter is larger (or smaller) than 10, the knot is deep (or shallow). YbeA has a chain length of 155 amino acids, and its knotted core extends from residue 69 to residue 121. Consequently, it is necessary to remove 68 residues from the N-terminus (or 34 residues from the C-terminus) to untie the knot that is classified as deep. The native structure of YibK also embeds a deep trefoil knot. Indeed, YibK has a chain length of 160 residues and a knotted core that extends from residue 77 through residue 120, making it necessary to remove 76 residues from the N-terminus (or 40 residues from the C-terminus) to untie the knot. Despite their structural similarity, these bacterial proteins share only 19% of sequence identity.
@@
To explore the role of the knotted topology in determining equilibrium properties, it is necessary to consider control systems Yeates2010. A suitable control system is a similar, yet unknotted protein that must be used for comparison. Since knotted/unknotted pairs do not exist in nature, they must be designed or engineered. In the present study, we consider two control systems obtained experimentally and one control system created through computational modeling. The in vitro control systems were engineered by Hsu and co-workers through a process known as circular permutation Danny2019a, Danny2019b. A circular permutant (CP) is an engineered protein that results from linking the C- and N-terminus of the polypeptide chain after disrupting the protein’s backbone at some selected site Faisca2012. In the case of protein YbeA, an unknotted control system was obtained by creating a CP after breaking the chain at the knotted core in posiiton 74, while for YibK the chain was broken at position 82, also in the knotted core, to create the CP (Figure 2A). In the present study, the resulting unknotted structures are designated by CP-YbeA (PDB ID: 5zyo) and CP-YibK (PDB ID: 6ahw). The in silico control structure was obtained for YibK (PDB ID: 1mxi) by manually passing the loop containing residues 114–125 across the knotted core (amino acids 81-85) (Figure 2B). The manipulated structure was then subjected to an MD simulation, in which segments 1-80, 86-113, and 126-156 were positionally restrained 1000 kJ/mol nm2), while the two loops (81–85 and 114–125) were allowed to relax. The final relaxed and refined structure is named IS-YibK.
@@
For all three structures, we computed the gyration radius (), which is a measure of structural compactness, the root-mean-square-deviation (RMSD) to the wild-type native structure, which is a measure of structural similarity, and the number of native contacts, which is a measure of the system’s energy in the context of the adopted model. These properties are reported in Table 1. The native structure of CP-YbeA exhibits significant structural rearrangements, which are visible in the cartoon representation (Figure 2A), and result in a substantially high RMSD (3Å). There is also a significant loss of compactness, resulting in a high gyration radius and a substantial loss (25%) of native contacts. On the other hand, both the unknotted CP-YibK and IS-YibK share a high degree of structural similarity with YibK (RMSD2Å). However, while IS-YibK has fewer native contacts than YibK (11%), the circular permutant exhibits a nearly identical degree of compactness, and essentially the same number of native contacts (the difference between the two model systems being only 2%), which makes it the best control system considered here.

### v2

# The role of topology on protein thermal stability

@@ Abstract
For several decades, experimental and computational studies have been used to investigate the potential functional role of knots in protein structures. A property that has attracted considerable attention is thermal stability, i.e., the extent to which a protein retains its native conformation and biological activity at high temperatures, without undergoing denaturation or aggregation. Thermal stability is quantified by the melting temperature , an equilibrium property that corresponds to the peak of heat capacity in differential scanning calorimetry (DSC) experiments. Experimental and computational studies report conflicting effects of knotting on protein thermal stability 40, 26. Here, we use extensive Monte Carlo simulations of a simple C-alpha model of protein YibK, with energetics modeled by the Go potential, to show that does not depend on the topological state of the protein. Our simulations further support the view that the discrepancy between the experimental and computational results stems from a pronounced separation of timescales for unknotting and unfolding that is inherent to deeply knotted proteins like YibK. In particular, the timescale separation implies that the complete unfolding–untying transition may not be accessible within the duration of a DSC experiment, whose apparent measurements likely reflect a non-equilibrium distribution lacking unfolded states that are also unknotted.
@@ 1 Introduction
A knotted protein is a protein whose polypeptide chain is self-tied, forming a physical (i.e., open) knot in its native structure. Although the first knotted protein was reported in 1977 39, it was not until 2000 that Taylor’s computational methodology 48 (which enables a systematic determination of the topological state of open polymer chains) brought these complex proteins into the spotlight. Knotted proteins make up 1% of all protein entries (225 k) in the Protein Data Bank (PDB) 7, and are statistically rare compared to compact random loops of the same size 28. All three domains of life contain knotted proteins 51, including Archaea, a group of single-celled organisms believed to have inhabited the planet for around 3.8 billion years. A recent survey that analyzed 700 k knotted proteins predicted by the computer artificial intelligence system AlphaFold (AF) 24, and 42 proteomes (16 bacteria, 9 fungi, 4 plants, and 13 animals), reported that the percentage of knotted proteins in any given proteome is around 0.4% 37.
The minimum number of crossings in all possible planar projections of a given chain is an incomplete topological invariant used to classify both topological knots (which are closed curves in space) and protein knots. The (or trefoil) knots, a class that features three crossings, is by far the most prevalent protein knot, followed by the and knots, with four and five crossings, respectively 7, 37. The most complex protein knot has seven crossings (it is a knot) and was predicted by AF 24. It is an 89-residue-long bacterial protein (termed Q9PR55) whose atomic structure was recently experimentally resolved, validating AF’s predictions 18.
@@
In the context of structural biology, the design principle ’form follows function’ means that a protein’s structure is adapted to successfully perform its specific task. It highlights the fact that protein structure is not arbitrary but is shaped by the specificities of its role, which ultimately contributes to the organism’s survival. The fact that knotted proteins are statistically rare does not imply a lack of function for the knot; rather, it indicates that folding into a knotted native structure is costly (e.g, it leads to slow folding and topological frustration 25) and therefore unlikely to be tolerated unless compensated by specific advantages. If knotting were purely neutral or deleterious, one would expect it to be entirely suppressed by evolution. The fact that knotted proteins have been conserved throughout evolution 45 (and, in many cases, have likely evolved from knotted precursor structures in organisms dating back billions of years 50) suggests that they may confer some advantage to their carriers. Thus, it is not surprising that a large body of theoretical and experimental work has been devoted to determining the potential functional roles of knots in proteins (reviewed in 21, 50), alongside establishing how they fold into their intricate native structures (reviewed in 13, 20, 19, 50). While significant advancements have been made regarding folding and knotting mechanisms, a consensus has not been reached concerning the function of protein knots 21. Knots can potentially influence proteins at both the level of physical properties, structure, as well as at the level of their biological function. A recent study found that the knot in protein UCHL-1 is critical for the preservation of secondary structure 15. Previous studies have reported that a knotted structure helps shape and form the binding site of enzymes 34, 35, 22. Additionally, one study has shown that the knotted region is directly involved in enzymatic activity 5. Indeed, for the protein TrmD (an example of an important class of enzymes, the SPOUT and SAM synthases), it was shown that the knot regulates catalysis, being critical to protein function 5. Examples of physical properties are thermal stability, which measures a protein’s resistance to melting and is quantified by the melting temperature; kinetic stability, which measures the protein’s resistance to unfolding at a certain temperature and is quantified by the unfolding rate constant; and mechanical stability, a measure of the protein’s ability to withstand physical forces and deformations without unfolding or losing its native structure. Simulation studies using models with different levels of resolution suggested that knots in proteins enhance kinetic stability 41, 46, while an enhancement of mechanical stability has been reported for several model systems based on in vitro experiments 43, 18, 19) and computational studies 46, 52.
Some years ago, we used Monte Carlo simulations of a simple lattice model to find that a knotted topology does not affect protein’s thermal stability 40. However, this contradicts a recent experimental study that reported a significant enhancement in the thermal stability of protein YibK, which features a deep trefoil knot 6. In lattice models, the discretization of conformational space restricts backbone geometry, hindering the study of specific proteins like YibK, which is only possible in the context of a continuum off-lattice model. Additionally, the reduced conformational freedom inherent to lattice models means that they cannot accurately capture entropic effects. Thus, the purpose of the present work is to revisit the question of whether or not thermal stability is affected by the presence of an open knot in the native structure of protein YibK. Our results strongly support the view that knots play no role in enhancing protein thermal stability or any other thermodynamic equilibrium property.
@@ 2.1 The Cα Gō model
Proteins are represented by a simple Cα model (Figure 1). Accordingly, residues are reduced to hard spherical beads of uniform size, centered on the Cα atoms. Consecutive Cα atoms are connected by rigid sticks representing pseudobonds on the amide planes. We adopt a radius of 1.7 Å for the beads, which is the van der Waals radius of Cα atoms 49. For the length of each stick, we adopt the distance between the Cα atoms of the respective bonded residues in the protein’s native conformation, these being approximately 2.9 Å for cis bonds and 3.8 Å for trans bonds. Two non-bonded residues are said to be in contact in the native conformation if the smallest distance between any two heavy atoms, one belonging to each residue, is 4.5 Å, this cut-off being chosen because it is slightly larger than twice the average van der Waals radius of heavy atoms in proteins.
To model protein energetics, we consider the native-centric Gō potential 16. Accordingly, the total energy of a conformation defined by bead coordinates is given by
@@ 2.2 Replica-exchange Monte Carlo simulations
The conformational space of the C-alpha model is explored with Metropolis Monte Carlo (MC) 31 by means of a move set that comprises crankshaft and pivot moves. All simulations start from an unfolded (and unknotted) conformation obtained from a high-temperature simulation, and folding progress is monitored using the fraction of native contacts, . To sample equilibrium distributions, we use Monte Carlo replica-exchange (MC-RE) 44.
To explore the conformational space, two sampling schemes can be deployed. A canonical one in which the move set preserves the linear topology of the chain, and another one that breaks the linear topology of the chain by allowing it to cross itself 9, while preserving the sequential bonding of the protein chain and excluded volume among its units. We have previously shown, using simple Cα protein models (of unknotted proteins Fn-III and m, shallowly knotted protein MJ0366, and deeply knotted protein Rds3p), that there is no difference in equilibrium properties calculated with a move set that preserves linear topology and one that does not 9. However, the move set that breaks the linear topology of the chain is extraordinarily advantageous in the case of deeply knotted proteins like the ones studied here, since it provides correct equilibrium results at a much lower computational cost 9. Thus, unless otherwise stated, we will deploy the sampling method that does not preserve the linear topology of the chain.
The weighted histogram analysis method (WHAM) 4 is used to analyze data from the MC-RE simulations and to produce maximum-likelihood estimates of the density of states, from which expected values for thermodynamic properties are calculated as functions of temperature. An example is the heat capacity, , defined in reduced units as . The melting temperature, , is determined as the temperature at which the peaks. We measure the cooperativity degree of the transition by the ratio of the full width at half maximum (FWHM) of the peak to the melting temperature, and the half-width of the potential well, , is adjusted to obtain a simulated FWHM ratio between 4 and 5% 27, 14. This criterion has been successfully used in previous simulations employing a similar potential and sampling 42, 10. WHAM is also used to project the density of states along to obtain free energy profiles at a specific temperature.
@@ 2.3 Molecular Dynamics simulations
The MD simulations were performed using GROMACS 2024.3 1, with the AMBER 14SB force fields 29, and the TIP3P water model 23, 33. The system consisted of a protein, 8808 water molecules, and 4 Cl- ions to equilibrate the overall system charges. We performed a two-step minimization procedure using the steepest descent method and an initialization protocol with two additional steps, in which temperature (NVT) and pressure (NPT) baths are turned on and equilibrated for 100 ps segments, with an integrator step of 1 fs. The nonbonded interactions were computed using a single cutoff of 1.0 nm, with the neighbor list updated every 20 steps. Van der Waals interactions were truncated beyond this cutoff, while long-range electrostatics were calculated using the Particle-Mesh Ewald (PME) method 8, with a Verlet cutoff of 1.0 nm and a Fourier grid spacing of 0.125 nm. Bond lengths involving hydrogen atoms in solute and water molecules were constrained using the LINCS 17 and SETTLE 32 algorithms, respectively. In the NVT ensemble, the v-rescale thermostat 2 was employed to maintain a temperature of 310 K (with a coupling constant of 0.1 ps). In the NPT ensemble, the v-rescale thermostat 2 (coupling constant of 1.0 ps) was used in combination with the Parrinello-Rahman barostat 36 (coupling constant of 2.0 ps and isothermal compressibility of 4.510-5 bar-1). For the MD relaxation step, we run for 100 ns with the integration step adjusted to 2 fs.
@@ 2.4 Topological state
The topological state (knotted or unknotted) of a sampled conformation is determined using the Koniaris-Muthukumar-Taylor (KMT) algorithm 48.
@@ 3.1 Model systems
Proteins YbeA (from Escherichia coli, PDB ID: 1ns5) and YibK (from Haemophilus influenzae, PDB ID: 1mxi and PDB ID:1j85) are homodimeric enzymes that belong to the \-knot superfamily of MTases. In this work, we consider the monomer identified as chain A in the corresponding PDB files. The native structure of both proteins features a trefoil knot. In a knotted protein, the region of the polypeptide chain encompassing the knot is termed the knotted core. Protein knots are classified as deep or shallow, depending on the number of residues that must be removed from one of the termini to unknot the chain. If the latter is larger (or smaller) than 20, the knot is deep (or shallow)47, 20, 38. YbeA has a chain length of 155 amino acids, and its knotted core extends from residue 69 to residue 121. Consequently, it is necessary to remove 68 residues from the N-terminus (or 34 residues from the C-terminus) to untie the knot that is classified as deep. The native structure of YibK also embeds a deep trefoil knot. Indeed, YibK has a chain length of 160 residues and a knotted core that extends from residue 77 to 120, requiring removal of 76 residues from the N-terminus (or 40 residues from the C-terminus) to untie the knot. Despite their structural similarity, these bacterial proteins share only 19% of sequence identity.
@@
To explore the role of the knotted topology in determining equilibrium properties, it is necessary to consider control systems 25. A suitable control system is a similar, yet unknotted protein that must be used for comparison. Since knotted/unknotted pairs do not exist in nature, they must be designed or engineered. In the present study, we consider two control systems obtained experimentally and one control system created through computational modeling. The in vitro control systems were engineered by Hsu and co-workers through a process known as circular permutation 6, 26. A circular permutant (CP) is an engineered protein that results from linking the C- and N-terminus of the polypeptide chain after disrupting the protein’s backbone at some selected site 12. In the case of protein YbeA, an unknotted control system was obtained by creating a CP after breaking the chain at the knotted core in position 74, while for YibK the chain was broken at position 82, also in the knotted core, to create the CP (Figure 2A). In the present study, the resulting unknotted structures are designated by CP-YbeA (PDB ID: 5zyo) and CP-YibK (PDB ID: 6ahw). The in silico control structure was obtained for YibK (PDB ID: 1mxi) by manually passing the loop containing residues 114–125 across the knotted core (amino acids 81-85) (Figure 2B). The manipulated structure was then subjected to an MD simulation, in which segments 1-80, 86-113, and 126-156 were positionally restrained 1000 kJ/mol nm2), while the two loops (81–85 and 114–125) were allowed to relax. The final relaxed and refined structure is named IS-YibK.
@@
For all structures, we computed the gyration radius (), which is a measure of structural compactness, the root-mean-square-deviation (RMSD) to the wild-type native structure, which is a measure of structural similarity, and the number of native contacts, which is a measure of the system’s energy in the context of the adopted model. These properties are reported in Table 1. The native structure of CP-YbeA exhibits significant structural rearrangements, which are visible in the cartoon representation (Figure 2A), and results in a substantially high RMSD (3Å). There is also a significant loss of compactness, resulting in a high gyration radius and a substantial loss (25%) of native contacts. On the other hand, both the unknotted CP-YibK and IS-YibK share a high degree of structural similarity with YibK (RMSD2Å). However, while IS-YibK has fewer native contacts than YibK (11%), the circular permutant exhibits a nearly identical degree of compactness and essentially the same number of native contacts (the difference between the two model systems being only 2%), which makes it the best control system considered here.